Asuming no heat loses heat released by condensation of xgof steam at 100C can be used to convert yg of ice at 0C into water at 100C ratio x:y is

Grade:12

1 Answers

Arun
5 years ago
 
according to your question,
x.Lv=y.Lf + y.100.s  where Lv and Lf are latent heat of vapourisation and fusion. and s is specific  =>  x.Lv=y(Lf + 100.s) heat of water
=>y/x = Lv/Lf + 100s, putting the values we get y/x = 3 :1
Hence x:y = 1: 3